OpenStudy (anonymous):

integrate sin^2n dx from 0 to pi

OpenStudy (anonymous):

is it sin^2x ?

OpenStudy (abb0t):

Do you mean: \[\int\limits_{0}^{2 \pi }\sin^2(x)dx\]

OpenStudy (abb0t):

sorry, i mean 0<x< pi

OpenStudy (anonymous):

i mean sin ^2n theta dtheta using cauch integral

OpenStudy (abb0t):

Using cauchy integral?! Really?!

OpenStudy (anonymous):

\[\int\limits_{0}^{\pi}\sin ^{2}ntheta \]

OpenStudy (anonymous):

yesss.. using cauchy integral. sin ^(2n)

OpenStudy (abb0t):

There's a formula for sin which You can use, I can't remember it at the top of my head, but it's something like: \[\frac{ e^x+e^{-ix} }{ 2 }\] I think? Does that look familiar to yopu?!

OpenStudy (abb0t):

For continuous functions, you use u(t) +iv(t) on a<theta<b?